# Asymptotic Schwarzschild solutions in $f(R)$ gravity and their observable effects on the photon sphere of black holes

### v1

@@ I Introduction
In this work, we analyze black hole solutions in $f(R)$ gravity that smoothly connect to the Schwarzschild solution in the exterior region. We focus on the model $f(R)=R+aR^{2}$, motivated by the assumption that the Ricci scalar remains small in the relevant regime delaCruz-Dombriz:2009pzc ; Nzioki:2009av . This model exhibits two distinct asymptotic behaviors: for $a>0$, the solutions decay exponentially, whereas for $a<0$, they show damped oscillations. Starting from the field equations, we obtain these asymptotic behaviors analytically and impose them as boundary conditions when integrating the system numerically with a fourth-order Runge–Kutta scheme. This guarantees that the resulting numerical solutions display the desired asymptotic properties.
@@ II.2 Results for $a<0$
We plot in Figure 4 results for $\alpha=1$ and $P_{\alpha}=0.1$. These solutions resemble the oscillatory behavior reported in AparicioResco:2016xcm . In this case, we can integrate from $x\gg 1$ down to the event horizon. An interesting feature emerges: these $f(R)$ black holes have a smaller horizon than the asymptotic SW mass would suggest. A distant observer infers a mass $M$ corresponding to $r_{S}$ (i.e., $x=1$), yet the actual horizon lies at $x<1$. Consequently, the photon-sphere radius is smaller and its width is modified. We now explore the dependence of the solution on $\alpha$ and $P_{\alpha}$. If we fix $P_{\alpha}$ and move $\alpha$ we find only modest differences for a range of $\alpha=0.05-1$. For lower values of $\alpha$ the oscillations have shorter period and amplitude but they do not change too much values of the metric $B(x)$. On the other hand, we can fix $\alpha$ and vary $P_{\alpha}$, results can be seen in Figure 6. In the limit $P_{\alpha}\to 0$ we recover SW solution but at bigger values of $P_{\alpha}$ we get lower radius for the event horizon.
@@ VI Conclusions
  • Finally, because our solutions are obtained numerically, it is essential to check whether they contain event or apparent horizons or lead to naked singularities. Modern horizon-finder algorithms, such as the hyperbolic–flow-based BHaHAHA library, recast the apparent-horizon equation into a damped nonlinear wave equation and achieve large speed-ups compared with traditional elliptic solvers Etienne2025 . Spectral algorithms can locate horizons with high accuracy on three-dimensional slices without requiring a root-finding step Lin2007 . Implementing these methods will clarify the causal structure of our solutions and ensure that the photon-sphere width is computed in a physically admissible spacetime.

### v3

@@ I Introduction
In this work, we analyze black hole solutions in $f(R)$ gravity that smoothly connect to the Schwarzschild solution in the exterior region. In a recent work, Scali and Piattella ScaliPiattella2024 construct a class of asymptotically Schwarzschild solutions in $f(R)$ gravity by perturbatively solving the modified Einstein equations and recovering the corresponding $f(R)$ a posteriori, which turns out to be non-analytic at $R=0$. However, we will focus on the model $f(R)=R+aR^{2}$ by assuming analytic $f(R)$ and that the Ricci scalar remains small in the relevant regime delaCruz-Dombriz:2009pzc ; Nzioki:2009av . We show that, in the limit of large radii, the Schwarzschild solution is asymptotically recovered. This model exhibits two distinct asymptotic behaviors: for $a>0$, the solutions decay exponentially, whereas for $a<0$, they show damped oscillations. Starting from the field equations, we obtain these asymptotic behaviors analytically and impose them as boundary conditions when integrating the system numerically with a fourth-order Runge–Kutta scheme. This guarantees that the resulting numerical solutions display the desired asymptotic properties.
@@ II.1 Results for $a>0$
It is important to emphasize that the numerical solutions obtained here should only be regarded as valid in the regime where the Ricci scalar remains small, that is, for radii not too close to the Ricci scalar divergence. As the curvature grows, the approximation $f(R)\simeq R+aR^{2}$ ceases to be reliable, and higher-order corrections in $R$ (such as $R^{3}$, $R^{4}$, etc.) would need to be included. Therefore, the divergent behavior of $R(r)$ and the metric functions at $x\ll 1$ is not to be interpreted as a physical prediction of the model but rather as an indication of the breakdown of the truncated expansion.
@@ II.2 Results for $a<0$
We plot in Figure 4 results for $\alpha=1$ and $P_{\alpha}=0.1$. These solutions resemble the oscillatory behavior reported in AparicioResco:2016xcm . In this case, we can integrate from $x\gg 1$ down to the event horizon ($B(x)\to 0$). An interesting feature emerges: these $f(R)$ black holes have a smaller horizon than the asymptotic SW mass would suggest. A distant observer infers a mass $M$ corresponding to $r_{S}$ (i.e., $x=1$), yet the actual horizon lies at $x<1$. Consequently, the photon-sphere radius is smaller and its width is modified. We now explore the dependence of the solution on $\alpha$ and $P_{\alpha}$. If we fix $P_{\alpha}$ and move $\alpha$ we find only modest differences for a range of $\alpha=0.05-1$. For lower values of $\alpha$ the oscillations have shorter period and amplitude but they do not change too much values of the metric $B(x)$. On the other hand, we can fix $\alpha$ and vary $P_{\alpha}$, results can be seen in Figure 6. In the limit $P_{\alpha}\to 0$ we recover SW solution but at bigger values of $P_{\alpha}$ we get lower radius for the event horizon.
@@
As shown in Figure 4, when $B(x)\to 0$ we find that $A(x)\to\infty$. From the metric in equation (4), it follows that the Ricci scalar remains finite at the effective event horizon. However, as $x\to 0$, the Ricci scalar grows rapidly and, as in the previous case, the approximation $f(R)\simeq R+aR^{2}$ is no longer valid. In this strong-curvature regime, higher-order corrections in $R$ would need to be taken into account to obtain physically meaningful results.
@@ VI Conclusions
  • Finally, we note that the truncated model $f(R)\simeq R+aR^{2}$ represents only the leading-order term of an analytic expansion around $R=0$. The pathologies appearing at small radii, such as the divergence of the Ricci scalar or the absence of a regular horizon, highlight the need to include higher-order terms in the curvature. Future work will therefore focus on extending the analysis to models of the form $f(R)=R+aR^{2}+bR^{3}+O(R^{4})$, in order to investigate whether such corrections can regularize the solutions and restore physically admissible black-hole metrics with proper horizons and singularities. For this purpose, modern horizon-finding algorithms Etienne2025 ; Lin2007 could be employed to accurately identify event horizons and curvature singularities in the numerical solutions. Such extensions would also allow for testing whether specific analytic forms of $f(R)$ can reconcile the existence of an event horizon with the asymptotic Schwarzschild behavior.